OpenStudy (anonymous):

Convert the following vectors; a. 7 N on a bearing of 145° to Cartesian form b. [-6, 8] to direction/magnitude form @Mertsj

OpenStudy (anonymous):

@Mertsj @Loser66 can you guys like heeelp??

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

step by step please :) @ybarrap

OpenStudy (ybarrap):

a) 7 N is the magnitude The magnitude in the x direction is $$ F_x=7 \times \cos 145 $$ The magnitude in the y direction is $$ F_y=7 \times \sin 145 $$ So the vector is $$ <F_x,F_y> $$

OpenStudy (anonymous):

Thanks so much !! and part b ? @ybarrap

OpenStudy (ybarrap):

b. [-6, 8] to direction/magnitude form $$ F=\sqrt{-6^2+8^2}\\ \theta =\arctan\cfrac{8}{-6} $$ Where magnitude is F and direction is \(\theta\)

OpenStudy (anonymous):

so that is the vector for b right? @ybarrap